OpenStudy (anonymous):

Factor by grouping 32x^2-8x-4

OpenStudy (andriod09):

\[32x^2-8x-4\] right?

OpenStudy (anonymous):

yep @andriod09

OpenStudy (andriod09):

okay, lets see. do you know how to factor in the first place?

OpenStudy (anonymous):

Yes, I have the answer,im just stuck as to how to get it. the asnwer is 4(4x+1)(2x-1)

OpenStudy (andriod09):

so then lets work from there. \[4(4x+1)(2x-1)\] what would you do first?

OpenStudy (anonymous):

FOIL

OpenStudy (andriod09):

do FOIL for me then, step by step, like do F, then do O, then I, then L in different posts, or use it slowly in the same posts. use line breaks, is what i mean.

OpenStudy (anonymous):

4x*2x=8x^2

OpenStudy (anonymous):

4x*1=4x

OpenStudy (anonymous):

1*2x=2x

OpenStudy (anonymous):

1*-1=-1 -------- O= should have been a -4 not 4.

OpenStudy (anonymous):

8x^2-4x+2x-1 -------------- 8x^2-2x-1

OpenStudy (andriod09):

you're missing something, redo your steps 1-3

OpenStudy (anonymous):

i thought you distribute the 4 at the end so then it would be 4(8x^2-2x-1)

OpenStudy (andriod09):

this is how i would do it, tell me if you see any mistakes. \[4(4x+1)(2x-1)\] \[(16x+4)(8x-4)\] then you would double it, as far as i know. i'm sorry if i didn't help. relized i haven't done this in nearly 3 years. xD

OpenStudy (anonymous):

Thaat sokay i figured out my missing step thanlks :D

OpenStudy (andriod09):

NP. thanks! and your welcome!